find the value of x
0=9(3x+7)-5(x+2)-(2x-5)

Answer by Cromlix(4381)

Hi there,
0=9(3x+7)-5(x+2)-(2x-5)
Multiply out
0 = 27x + 63 - 5x - 10 - 2x + 5
Collect like terms:
0 = 27x - 5x - 2x + 63 - 10 + 5
0 = 20x + 58
20x = -58
x = -58/20
x = -29/10 = -2.9
Hope this helps :-)